# Geodesic Slice Sampler for Multimodal Distributions with Strong Curvature

### v1

@@ Abstract
Traditional Markov Chain Monte Carlo sampling methods often struggle with sharp curvatures, intricate geometries, and multimodal distributions. Slice sampling can resolve local exploration inefficiency issues and Riemannian geometries help with sharp curvatures. Recent extensions enable slice sampling on Riemannian manifolds, but they are restricted to cases where geodesics are available in closed form. We propose a method that generalizes Hit-and-Run slice sampling to more general geometries tailored to the target distribution, by approximating geodesics as solutions to differential equations. Our approach enables exploration of regions with strong curvature and rapid transitions between modes in multimodal distributions. We demonstrate the advantages of the approach over challenging sampling problems.
@@ 1 Introduction
Sampling from a differentiable unnormalized log-density defined on a Euclidean space is a core problem in machine learning and statistics. While gradient-based Markov Chain Monte Carlo (MCMC) methods have proven effective in many scenarios, they often face significant challenges when the target distribution exhibits complex geometry (sharp curvature) or multimodal behavior. The two core challenges are largely addressed with complementary techniques, with little work in algorithms that excel for targets that are both multimodal and complex in shape.
@@
Complex shapes and sharp curvatures are often addressed by using a suitably chosen Riemannian geometry within the sampling algorithms [Girolami and Calderhead, 2011]. Instead of operating in a Euclidean space and metric, the samplers carry out the necessary operations using a metric that adapts to the curvature of the parameter space. In practice, the methods follow flows induced by the metric, in most cases by numerical integration, and consequently the methods are sometimes called geodesic methods as in our title. Various practical metrics and sampling algorithms have been shown to improve sampling of targets with strong curvature [Girolami and Calderhead, 2011, Byrne and Girolami, 2013, Lan et al., 2015, Hartmann et al., 2022, Yu et al., 2024, Williams et al., 2024], albeit always with increased computational cost.
@@
Multimodality, in turn, is most commonly addressed by tempering or diffusion techniques [Earl and Deem, 2005, Chen et al., 2024]. These methods use tempered – smoothed – version of the target to improve exploration over multiple modes, intuitively changing the problem itself so that the modes are connected with areas of sufficient probability. At high degree of tempering these methods can efficiently explore the different modes, but low tempering is needed for accurate sampling within the modes, necessitating adaptive or parallel sampling with different degrees of tempering. The efficiency of parallel tempering depends on the swap acceptance rate between adjacent temperatures, which can decrease in high dimensions if the temperature schedule is not well-tuned [Woodard et al., 2009]. Diffusion-based approaches, in turn, require careful choice of the noise schedule to balance exploration and accuracy [Song and Ermon, 2019, Chen et al., 2024]. Unlike tempering, diffusion methods can achieve smooth transitions between modes without explicitly maintaining a set of parallel chains, but the acceptance rate of noisy samples can be low [Chen et al., 2024].
@@
Even though the two approaches are efficient in addressing the two challenges separately, there is very little work on samplers designed for the general setup where both difficulties may arise simultaneously. One could consider e.g. parallel tempering in a Riemannian metric – see Byrne and Girolami [2013] for a rare example in this intersection – but ideally we would like to address both aspects using a common mechanism. This work explores one such approach, developing a Riemannian sampler capable of efficiently exploring multiple modalities, without any tempering for the target distribution. Instead, we seek to improve mode exploration by change of the metric, in the spirit of the early work by Lan et al. [2014] that developed a specific metric solely for this purpose. Their metric, however, requires explicit identification and tracking of the modes and is more like a conceptual demonstration, and we are not aware of any other works aiming for efficient multi-modal samplers solely by the change of the metric.
@@
Our starting point is the idealized slice sampler, with computable level sets. The performance of such an idealized sampler is theoretically independent of the sampling problem, as pointed out by Durmus et al. [2023]: “This means that the performance of the idealized slice sampler is ignorant of the introduction of, e.g., multimodality, local modes, or anisotropy as long as the volume of the level sets is not modified.” From a practical perspective, we build on the (Euclidean) Hit-and-Run slice sampler by Bélisle et al. [1993], which at each iteration selects a random direction and then samples from the resulting one-dimensional distribution formed by the intersection of the line and the slice. In effect, it transforms multi-dimensional sampling into sequential one-dimensional sampling tasks, but the overall sampler may be inefficient. Especially in higher dimensions the intersection with the slice can be small for almost all directions Murray et al. [2010].
@@
Both Habeck et al. [2023] and Durmus et al. [2023] recently considered generalizations of the Hit-and-Run sampler for Riemannian manifolds, replacing the lines with geodesics. We build on the general algorithmic framework introduced by Durmus et al. [2023] and adapt it to the task of sampling from a distribution with a complex geometry. Specifically, we begin by embedding the (Euclidean) sampling space into a higher-dimensional space that incorporates the target distribution’s geometric information, such as Fisher information or Monge embedding [Hartmann et al., 2022]. This transforms the problem into sampling from a particular Riemannian manifold where the target distribution corresponds to the Hausdorff density (see Section 3). Note that even though we leverage components proposed by Durmus et al. [2023], our task is fundamentally more difficult. Their starting point was sampling of a density on a known manifold (e.g. a sphere) where the geodesics are known exactly, whereas the complexity of our embedding manifold requires us to approximate the geodesics using numerical integrators.
@@
In this work, we propose a geodesic slice sampler applicable for arbitrary Riemannian metrics, and discuss the choice of the metric. In particular, we introduce two new computationally efficient metrics that both improve sampling over multimodal targets by, in a sense, pulling the modes closer to each other; see Figure 1 illustrating this effect within the slice sampler, as a function of a parameter $\lambda$ controlling how much the metric warps the space. In addition, we introduce a meta-sampler similar to Tjelmeland and Hegstad [2001] that combines the proposed method with separate sampler for improved exploration of local modes.
We empirically demonstrate improved sampling over Euclidean methods for complex targets, and highlight improved mixing over multiple modes in high dimensional-cases when compared against parallel tempering [Swendsen and Wang, 1986, Łatuszyński et al., 2025] and diffusive Gibbs sampler by Chen et al. [2024] designed for addressing multimodality. Similar to previous Riemannian methods, the algorithm shows good exploration and mixing but has slower iterations, because of computation of the geodesics.
@@ 2 Background: Slice sampling
To address this, the step-out and shrinkage procedures are used. Below, we provide an informal explanation of these procedures. The full algorithm is detailed in the appendix (Algorithms 3 and 4). Both procedures were first introduced by Neal [2003], but we adopt a modified version of the shrinkage step as proposed by Durmus et al. [2023]. Both versions are equally valid. For a moment, assume a univariate density $p(x)$ and a current position $x\in\operatorname{\mathbb{R}}$. The procedures are as follows:
@@
Figure 2: The Hit-and-Run procedure on the Funnel distribution with base point ()($\star$)( ⋆ ). The green circle is $\mathbb{S}^{D-1}$, and the red arrow is a norm one velocity. On the left the space is Euclidean and the straight line scapes the distribution while the right for the Monge metric $\alpha=1$ the geodesic follows curves towards the narrow area.
@@
One way to extend slice sampling to multivariate distributions is to combine it with Hit-and-Run sampling, here presented following Bélisle et al. [1993]. Let $\operatorname{\mathbb{S}}^{D-1}(\operatorname{\boldsymbol{x}})=\{\operatorname{\boldsymbol{v}}\in\operatorname{\mathbb{R}}^{D}:\norm{\operatorname{\boldsymbol{v}}}^{2}=1\}$, and $\operatorname{\boldsymbol{v}}\sim\mathrm{Unif}(\operatorname{\mathbb{S}}^{D-1}(\operatorname{\boldsymbol{x}}))$. An iteration of the whole sampler is:
@@ 3 Method
Straight-line Hit-and-Run sampling can be inefficient because proposals often move away from high-probability regions [Murray et al., 2010]. To resolve this, we perform slice sampling along geodesic curves that can accommodate the geometry of the target distribution. This improves efficiency when the target distribution is highly curved, or multimodal distributions; see Section 4.2. We are interested in sampling problems defined in $\mathbb{R}^{d}$ but allow using different plug-in metrics (preferably using the target density information) to enhance exploration. This can be cast as sampling from a distribution defined on a Riemannian manifold where the algorithm in Durmus et al. [2023] can be employed. Alternatively, it can be seen as Hit-and-Run where straight lines are replaced by curves that better wrap around the level sets of the target density (given the metrics are good enough). However, since the metric is quite general, geodesics can not be obtained in closed-forms and we need to use numerical integrators. To correctly sample along geodesics, we need three key components: Adjusting for the correct density on the manifold, properly sampling directions using the Riemannian metric, and solving the geodesic equations.
@@ 3.1 Meta Sampler and Multimodality
The sampler as described above is valid as such, but we also introduce a simple extension that can further improve sampling for multi-modal targets with complex local structure.
@@ 4 Metrics
The sampler is general, applicable for arbitrary metric and only requiring $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ to be positive definite and vary continuously. By selecting an appropriate metric we can influence how the geodesics explore the space, controlling the overall sampling behavior. There is no single metric that is optimal for all targets, and the metrics proposed in the literature are motivated by complementary argumentation, with notable emphasis in computational efficiency.
@@ 4.1 For adapting to local curvature
The Fisher Information Metric (FIM) is defined as the covariance of the score function, and was predominantly used in the early Riemannian methods [Girolami and Calderhead, 2011] due to its close connection to estimation theory. A general form of the metric is
@@
$\operatorname{\boldsymbol{G}}_{M}^{-1}(\operatorname{\boldsymbol{x}})$ $=\operatorname{\boldsymbol{I}}_{D}-\frac{\alpha^{2}}{1+\alpha^{2}\|\nabla\ell\|^{2}}\nabla\ell\nabla\ell^{\top}.$ (3)
As $\alpha^{2}\to 0$, the metric reduces to the Euclidean metric $\operatorname{\boldsymbol{I}}_{D}$. The determinant required for computing the Hausdorff density (Equation (1)) is $\det\operatorname{\boldsymbol{G}}_{M}(\operatorname{\boldsymbol{x}})=1+\alpha^{2}\|\nabla\ell\|^{2}$. Figure 3 illustrates the exponential map of geodesic balls with increasing radius under the Monge metric. This metric adapts to the geometry of the target distribution, expanding regions based on the local structure of the density.
@@
Another efficient metric, the Generative metric that is proportional to the target density function, was recently proposed by [Kim et al., 2024]. One of its advantages is that computing the Christoffel symbols $\Gamma^{k}_{ij}$ only requires first-order derivatives of the density, whereas the Monge metric (Equation 4.1) introduces second-order terms. For $p_{0}>0$ and $\lambda\geq 0$, the Generative metric and its inverse are:
$\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p(\operatorname{\boldsymbol{x}})+\lambda}{p_{0}+\lambda}\right)^{2}\boldsymbol{I}_{D},$ (4)
$\operatorname{\boldsymbol{G}}_{g}^{-1}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}})+\lambda}\right)^{2}\boldsymbol{I}_{D}.$ (5)
As $\lambda\to\infty$, the metric reduces to the Euclidean metric. The determinant is $\det\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})=\left(\tfrac{p(\operatorname{\boldsymbol{x}})+\lambda}{p_{0}+\lambda}\right)^{2D}$. Figure 1 illustrates the effect of $\lambda$ on the Hausdorff density along geodesics $t\mapsto p_{\mathcal{H}}\left(\hat{\gamma}_{(\operatorname{\boldsymbol{x}},\operatorname{\boldsymbol{v}})}(t)\right)$, and Figure 3 again shows how the metric transforms the space.
@@ 4.2 For Bridging the Modes
The above metrics adapt for the local curvature and have been designed to improve sampling of, for instance, narrow funnels by re-defining the proximity (see Figure 3). For assisting exploration of multimodal targets we need different kinds of metrics: Now we would want a metric that makes modes that are far away in the original Euclidean sense appear closer (see Figure 1). With the exception of the construction of Lan et al. [2014], which we will discuss more closely in Section 6, we are not aware of any previous metrics designed for this. Next, we introduce two computationally efficient metrics, with fast inverses and determinants, for assisting multi-modal sampling.
We start by reminding that any matrix $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ defines a valid Riemannian metric as long as it is positive definite for every $\operatorname{\boldsymbol{x}}\in\operatorname{\mathcal{M}}$ and varies continuously on $\operatorname{\mathcal{M}}$. Since the inverse of a positive definite matrix is also positive definite, we observe that it is possible to use any of the previously formulated $\operatorname{\boldsymbol{G}}^{-1}(\operatorname{\boldsymbol{x}})$ for defining a metric. Concretely, we propose:
@@
$\operatorname{\boldsymbol{G}}_{IM}(\operatorname{\boldsymbol{x}})$ $=\operatorname{\boldsymbol{I}}_{D}-\frac{\alpha^{2}}{1+\alpha^{2}\|\nabla\ell\|^{2}}\nabla\ell\nabla\ell^{\top}.$
as the metric tensor, with the inverse $\operatorname{\boldsymbol{G}}_{IM}^{-1}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{G}}_{M}(\operatorname{\boldsymbol{x}})$ given by the previously introduced metric tensor of the standard Monge metric (Eq. (4.1)). The determinant of this metric is $1/\det(\operatorname{\boldsymbol{G}}_{M})$, and hence it retains the computational efficiency of the original Monge metric.
@@
Similarly, we define
$\operatorname{\boldsymbol{G}}_{Ig}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}})+\lambda}\right)^{2}\boldsymbol{I}_{D},$
and obtain the inverse $\operatorname{\boldsymbol{G}}_{Ig}^{-1}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})$ as the metric tensor of the standard Generative metric (Eq. (4)) and the determinant as $1/\det(\operatorname{\boldsymbol{G}}_{g})$. Again, the computational efficiency of the original Generative metric is retained.
Next, we explain how the inverse Monge and inverse Generative achieve the desired mode-bridging ability. Geodesic curves maintain a constant velocity norm in the Riemannian sense by construction. Let $\gamma_{(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})}(t)=\operatorname{\boldsymbol{x}}_{t}$ be a geodesic curve with velocity $\dot{\gamma}_{(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})}(t)=\operatorname{\boldsymbol{v}}_{t}$, starting from $\operatorname{\boldsymbol{x}}_{0}$ with initial velocity $\operatorname{\boldsymbol{v}}_{0}$. If the geodesic moves toward a low-probability region where $p(\operatorname{\boldsymbol{x}}_{t})\to 0$, then we the “mode bridging” behavior occurs if the Euclidean velocity norms satisfy $\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}\gg\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$. This means that as $t$ increases in low-density regions, geodesics traverse these areas quickly, bringing distant modes closer. Figure 1 illustrates this effect for the inverse Generative metric. And the following observation collects this idea which we discuss further in Appendix A.6:
Observation 1.
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ with respect to the inverse Generative/Monge metric. Then, for $t$ such that $p(\operatorname{\boldsymbol{x}}_{t})\to 0$, the Euclidean velocity norm $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$ gets small, implying the low value density regions are covered faster as function of time $t$.

### v2

@@ Abstract
Traditional Markov Chain Monte Carlo sampling methods often struggle with sharp curvatures, intricate geometries, and multimodal distributions. Slice sampling can resolve local exploration inefficiency issues, and Riemannian geometries help with sharp curvatures. Recent extensions enable slice sampling on Riemannian manifolds, but they are restricted to cases where geodesics are available in a closed form. We propose a method that generalizes Hit-and-Run slice sampling to more general geometries tailored to the target distribution, by approximating geodesics as solutions to differential equations. Our approach enables the exploration of the regions with strong curvature and rapid transitions between modes in multimodal distributions. We demonstrate the advantages of the approach over challenging sampling problems.
@@ 1 Introduction
Sampling from a differentiable unnormalized log-density defined on a Euclidean space is a core problem in machine learning and statistics. While gradient-based Markov Chain Monte Carlo (MCMC) methods have proven effective in many scenarios, they often face significant challenges when the target distribution exhibits complex geometry (sharp curvature) or multimodal behavior. The two core challenges are largely addressed with complementary techniques, with little work on algorithms that excel for targets that are both multimodal and complex in shape.
@@
Complex shapes and sharp curvatures are often addressed by using a suitably chosen Riemannian geometry within the sampling algorithms [Girolami and Calderhead, 2011]. Instead of operating in a Euclidean space and metric, the samplers carry out the necessary operations using a metric that adapts to the curvature of the parameter space. In practice, the methods follow flows induced by the metric, in most cases by numerical integration, and consequently the methods are sometimes called geodesic methods as in our title. Various practical metrics and sampling algorithms have been shown to improve the sampling of targets with strong curvature [Girolami and Calderhead, 2011, Byrne and Girolami, 2013, Lan et al., 2015, Hartmann et al., 2022, 2023, Williams et al., 2024], albeit always with increased computational cost.
@@
Multimodality, in turn, is most commonly addressed by tempering or diffusion techniques [Earl and Deem, 2005, Chen et al., 2024]. These methods use a tempered (smoothed) version of the target to improve exploration over multiple modes, intuitively changing the problem itself so that the modes are connected with areas of sufficient probability. At a high degree of tempering these methods can efficiently explore the different modes, but low tempering is needed for accurate sampling within the modes, necessitating adaptive or parallel sampling with different degrees of tempering. The efficiency of parallel tempering depends on the swap acceptance rate between adjacent temperatures, which can decrease in high dimensions if the temperature schedule is not well-tuned [Woodard et al., 2009]. Diffusion-based approaches, in turn, require careful choice of the noise schedule to balance exploration and accuracy [Song and Ermon, 2019, Chen et al., 2024]. Unlike tempering, diffusion methods can achieve smooth transitions between modes without explicitly maintaining a set of parallel chains, but the acceptance rate of noisy samples can be low [Chen et al., 2024].
@@
Even though the two approaches are efficient in addressing the two challenges separately, there is very little work on samplers designed for the general setup where both difficulties may arise simultaneously. One could consider e.g. parallel tempering in a Riemannian metric — see Byrne and Girolami [2013] for a rare example in this intersection — but ideally we would like to address both aspects using a common mechanism. This work explores one such approach, developing a Riemannian sampler capable of efficiently exploring multiple modes, without any tempering for the target distribution. Instead, we seek to improve mode exploration by changing the metric, in the spirit of the early work by Lan et al. [2014] that developed a specific metric solely for this purpose. Their metric, however, requires explicit identification and tracking of the modes and is more like a conceptual demonstration, and we are not aware of any other works aiming for efficient multimodal samplers solely by the change of the metric.
@@
We are motivated by the idealized slice sampler with computable level sets. As noted by Durmus et al. [2023]: “This means that the performance of the idealized slice sampler is ignorant of the introduction of, e.g., multimodality, local modes, or anisotropy as long as the volume of the level sets is not modified.” This insight suggests that by modifying the geometry of the problem to produce simpler or more tractable level sets, the slice sampler can effectively handle multimodal distributions. From a practical perspective, we build on the (Euclidean) Hit-and-Run slice sampler by Bélisle et al. [1993], which at each iteration selects a random direction and then samples from the resulting one-dimensional distribution formed by the intersection of the line and the slice. In effect, it transforms multi-dimensional sampling into sequential one-dimensional sampling tasks, but the overall sampler may be inefficient. Especially in higher dimensions, the intersection with the slice can be small for almost all directions Murray et al. [2010].
@@
Both Habeck et al. [2023] and Durmus et al. [2023] recently considered generalizations of the Hit-and-Run sampler for Riemannian manifolds, replacing the lines with geodesics. We build on the general algorithmic framework introduced by Durmus et al. [2023] and adapt it to the task of sampling from a distribution with a complex geometry. Specifically, we begin by embedding the (Euclidean) sampling space into a higher-dimensional space that incorporates the target distribution’s geometric information, such as Fisher information or Monge embedding [Hartmann et al., 2022]. This transforms the problem into sampling from a particular Riemannian manifold where the target distribution corresponds to the Hausdorff density (see Section 3). Note that even though we leverage components proposed by Durmus et al. [2023], our task is fundamentally more difficult. Their starting point was sampling of a density on a known manifold (e.g., a sphere) where the geodesics are exactly known, whereas the complexity of our embedding manifold requires us to approximate the geodesics using numerical integrators.
@@
In this work, we propose a geodesic slice sampler applicable for arbitrary Riemannian metrics, and discuss the choice of the metric. In particular, we introduce two new computationally efficient metrics. Both metrics improve sampling over multimodal targets by, in a sense, pulling the modes closer to each other; see Figure 1 illustrating this effect within the slice sampler, as a function of a parameter $\lambda$ controlling how much the metric warps the space. In addition, we introduce a meta-sampler similar to Tjelmeland and Hegstad [2001] that combines the proposed method with a separate sampler for improved exploration of local modes.
We empirically demonstrate improved sampling over Euclidean methods for complex targets, and highlight improved mixing over multiple modes in high dimensional-cases when compared against parallel tempering [Swendsen and Wang, 1986, Łatuszyński et al., 2025] and the diffusive Gibbs sampler by Chen et al. [2024] designed for addressing multimodality. Similar to previous Riemannian methods, the algorithm shows good exploration and mixing, but has slower iterations because of the numerical computation of the geodesics.
@@ 2 Background: Slice sampling
To address this, the step-out and shrinkage procedures are used. Below, we provide an informal explanation of these procedures. The full algorithm is detailed in the Appendix (Algorithms 3 and 4). Both procedures were first introduced by Neal [2003], but we adopt an equally valid modified version of the shrinkage step as proposed by Durmus et al. [2023]. For a moment, assume a univariate density $p(x)$ and a current position $x\in\operatorname{\mathbb{R}}$. The procedures are as follows:
@@
One way to extend slice sampling to multivariate distributions is to combine it with Hit-and-Run sampling, presented here following Bélisle et al. [1993]. Let $\operatorname{\mathbb{S}}^{D-1}(\operatorname{\boldsymbol{x}})=\{\operatorname{\boldsymbol{v}}\in\operatorname{\mathbb{R}}^{D}:\norm{\operatorname{\boldsymbol{v}}}^{2}=1\}$, and $\operatorname{\boldsymbol{v}}\sim\mathrm{Unif}(\operatorname{\mathbb{S}}^{D-1}(\operatorname{\boldsymbol{x}}))$. An iteration of the whole sampler is:
@@
  • Obtain a sample from the density evaluated on the straight line (Euclidean geodesic)
@@ 3 Method
Straight-line Hit-and-Run sampling can be inefficient because proposals often move away from high-probability regions [Murray et al., 2010]. To resolve this, we perform slice sampling along geodesic curves that can accommodate the geometry of the target distribution. This improves efficiency when the target distribution is highly curved or multimodal; see Section 4.2. We are interested in sampling problems defined in $\mathbb{R}^{D}$ but allow using different plug-in metrics (preferably using the target density information) to enhance exploration.
The general problem can be cast as sampling from a distribution defined on a Riemannian manifold where we adapt the algorithm of Durmus et al. [2023] under general metrics. Alternatively, it can be seen as Hit-and-Run where straight lines are replaced by curves that better wrap around the level sets of the target density (given the metrics are good enough). Because the metric is general, closed-form geodesics are unavailable, so we must compute them with numerical integrators. To correctly sample along geodesics, we need three key components: Adjusting for the correct density on the manifold, properly sampling directions using the Riemannian metric, and solving the geodesic equations.
@@ 3.1 Meta Sampler and Multimodality
The sampler as described above is valid as such, but we also introduce a simple extension that can further improve sampling for multimodal targets with complex local structure.
@@ 4 Metrics
The sampler is general, applicable for an arbitrary metric and only requiring $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ to be positive definite and vary continuously. By selecting an appropriate metric we can influence how the geodesics explore the space, controlling the overall sampling behavior. There is no single metric that is optimal for all targets, and the metrics proposed in the literature are motivated by complementary argumentation, with notable emphasis in computational efficiency.
@@ 4.1 For adapting to local curvature
The Fisher Information Metric (FIM) is defined as the covariance of the score function, and was predominantly used in the early Riemannian methods [Girolami and Calderhead, 2011] due to its close connection to estimation theory. A general form of the metric is:
@@
$\operatorname{\boldsymbol{G}}_{M}^{-1}(\operatorname{\boldsymbol{x}})$ $=\operatorname{\boldsymbol{I}}_{D}-\frac{\alpha^{2}}{1+\alpha^{2}\|\nabla\ell\|^{2}}\nabla\ell\nabla\ell^{\top},$ (3)
where $\ell(x)=\ln p(x)$. As $\alpha^{2}\to 0$, the metric reduces to the Euclidean metric $\operatorname{\boldsymbol{I}}_{D}$. The determinant required for computing the Hausdorff density (Eq. (1)) is $\det\operatorname{\boldsymbol{G}}_{M}(\operatorname{\boldsymbol{x}})=1+\alpha^{2}\|\nabla\ell\|^{2}$. Figure 2 illustrates the exponential map of geodesic balls with increasing radius under the Monge metric. This metric adapts to the geometry of the target distribution, expanding regions based on the local structure of the density.
@@
Another efficient metric, the Generative metric that is proportional to the target density function, was recently proposed by Kim et al. [2024]. One of its advantages is that computing the Christoffel symbols $\Gamma^{k}_{ij}$ only requires first-order derivatives of the density, whereas the Monge metric (Equation 4.1) introduces second-order terms. For scalars $p_{0}>0$ and $\lambda\geq 0$, the Generative metric and its inverse are:
$\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}})+\lambda}\right)^{2}\boldsymbol{I}_{D},$ (4)
$\operatorname{\boldsymbol{G}}_{g}^{-1}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p(\operatorname{\boldsymbol{x}})+\lambda}{p_{0}+\lambda}\right)^{2}\boldsymbol{I}_{D}.$ (5)
As $\lambda\to\infty$, the metric reduces to the Euclidean metric. The determinant is $\det\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})=\left(\tfrac{p_{0}+\lambda}{p(\operatorname{\boldsymbol{x}})+\lambda}\right)^{2D}$. Figure 1 illustrates the effect of $\lambda$ on the Hausdorff density along geodesics $t\mapsto p_{\mathcal{H}}\left(\hat{\gamma}_{(\operatorname{\boldsymbol{x}},\operatorname{\boldsymbol{v}})}(t)\right)$, and Figure 2 again shows how the Generative metric transforms the space.
@@ 4.2 For Bridging the Modes
The above metrics adapt for the local curvature and have been designed to improve sampling of, for instance, narrow funnels by re-defining the proximity (see Figure 2). For assisting exploration of multimodal targets we need different kinds of metrics: Now we would want a metric that makes modes that are far away in the original Euclidean sense appear closer. With the exception of the construction of Lan et al. [2014], which we will discuss in Section 6, we are not aware of any previous metrics designed for this. Next, we introduce two computationally efficient metrics, with fast inverses and determinants, for assisting multimodal sampling.
Geodesic curves maintain a constant velocity norm in the Riemannian sense by construction. Let $\operatorname{\boldsymbol{x}}_{t}=\gamma_{(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})}(t)$ be a geodesic curve with velocity $\operatorname{\boldsymbol{v}}_{t}=\dot{\gamma}_{(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})}(t)$, starting from $\operatorname{\boldsymbol{x}}_{0}$ with initial velocity $\operatorname{\boldsymbol{v}}_{0}$. If the geodesic moves toward a low-probability region where $p(\operatorname{\boldsymbol{x}}_{t})\to 0$, then the “mode bridging” behavior occurs if the Euclidean velocity norms satisfy $\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}\ll\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}$. This means that as $t$ increases in low-density regions, the geodesics curves accelerate and locally pull the distant modes closer.
We propose two metrics with the desired behavior, by leveraging the metrics described in Section 4.1 in a novel way. Any matrix $\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})$ defines a valid Riemannian metric as long as it is positive definite for every $\operatorname{\boldsymbol{x}}\in\operatorname{\mathcal{M}}$ and varies continuously on $\operatorname{\mathcal{M}}$. Since the inverse of a positive definite matrix is also positive definite, we observe that it is possible to use any of the previously formulated $\operatorname{\boldsymbol{G}}^{-1}(\operatorname{\boldsymbol{x}})$ for defining a metric. This gives two new metrics that both help exploring multiple modes in different ways:
@@
$\operatorname{\boldsymbol{G}}_{IM}(\operatorname{\boldsymbol{x}})$ $=\operatorname{\boldsymbol{I}}_{D}-\frac{\alpha^{2}}{1+\alpha^{2}\|\nabla\ell\|^{2}}\nabla\ell\nabla\ell^{\top},$
as the metric tensor, with the inverse $\operatorname{\boldsymbol{G}}_{IM}^{-1}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{G}}_{M}(\operatorname{\boldsymbol{x}})$ given by the previously introduced metric tensor of the standard Monge metric (Eq. (4.1)). The determinant of this metric is $1/\det(\operatorname{\boldsymbol{G}}_{M})$, and hence it retains the computational efficiency of the original Monge metric. Figure 3 illustrates the geodesics emanating from one mode of a bimodal distribution under the Inverse Monge metric. The metric twists the curves towards the second mode and slightly increases acceleration (seen by the change of color). Observation 1 mathematically states the conditions for the change of acceleration caused by the metric.
Observation 1.
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ with respect to the Inverse Monge metric, such that $\operatorname{\boldsymbol{x}}_{0}$ is a local maximum. Then $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}\geq\norm{\operatorname{\boldsymbol{v}}_{0}}_{2}$ for all $t\neq 0$.
@@
We use
$\operatorname{\boldsymbol{G}}_{Ig}(\operatorname{\boldsymbol{x}})$ $=\left(\frac{p(\operatorname{\boldsymbol{x}})+\lambda}{p_{0}+\lambda}\right)^{2}\boldsymbol{I}_{D},$
and obtain the inverse $\operatorname{\boldsymbol{G}}_{Ig}^{-1}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{G}}_{g}(\operatorname{\boldsymbol{x}})$ as the metric tensor of the standard Generative metric (Eq. (4)) and the determinant as $1/\det(\operatorname{\boldsymbol{G}}_{g})$. Again, the computational efficiency of the original Generative metric is retained. Figure 3 illustrates the main effect of the metric, that is to accelerate on low density regions (indicated by the light color); it also twists the trajectories slightly (best seen within the initial mode and beyond the second mode in the top right corner). Additionally Figure 1 illustrates the behavior in a univariate distribution. The acceleration behavior is mathematically stated in Observation 2.
Observation 2.
Let $p(\operatorname{\boldsymbol{x}})$ be a smooth density function. Let $(\operatorname{\boldsymbol{x}}_{t},\operatorname{\boldsymbol{v}}_{t})$ be the geodesic flow with initial conditions $(\operatorname{\boldsymbol{x}}_{0},\operatorname{\boldsymbol{v}}_{0})$ such that $p(x_{0})>0$ with respect to the Inverse Generative metric. Then, for $t$ such that $p(\operatorname{\boldsymbol{x}}_{t})\to 0$ we have $\norm{\operatorname{\boldsymbol{v}}_{t}}_{2}>\norm{\operatorname{\boldsymbol{v}}_{t}}_{0}$.
The mathematical details for Observations 1 and 2 are given in Appendix A.6.
Figure 3: Effect of the metric for multimodal targets, showing the geodesics (lines) and the relative compression of the distance (color; yellow means faster travel in that area, darker colors mean slower travel). Left: Inverse Monge metric ($\alpha=0.001$) helps more the geodesics to reach the other mode, and also slightly compresses the distances in the low-probability region. Right: Inverse Generative metric ($\lambda=1$) compresses the distances in the low-probability region, but twists the paths only slightly.